OpenStudy (anonymous):

how do you simplify this ... x^3+1

OpenStudy (anonymous):

you cannot "simplify" it you can, however, factor it

OpenStudy (theviper):

\(\LARGE{x^3+1=??}\)

OpenStudy (anonymous):

you can use \(a^3+b^3=(a+b)(a^2-ab+b^2)\) with \(a=x,b=1\)

OpenStudy (anonymous):

Do you mean factorise x^3 +1 = 0?

OpenStudy (anonymous):

but there is no way that the factored form is any "simpler" than \(x^3-1\)

OpenStudy (anonymous):

or even \(x^3+1\) it is just a different form

OpenStudy (anonymous):

can you expand it??

OpenStudy (anonymous):

it is already expanded

OpenStudy (anonymous):

standard form is \(x^3+1\) factored form is \((x+1)(x^2-x+1)\)

OpenStudy (anonymous):

@Lulu212 perhaps you are getting this confused with \((x+1)^3\) which you can expand